what is the answer to logx + log (x+15) = 2?
Answer by user_dude2008(1825)

log is base 10

log(x) + log (x+15) = 2


log(x(x+15)) = 2

10^2=x^2+15x

x^2+15x-100=0

(x+20)(x-5)=0

x=-20 (extraneous) or x=5


Ans: only solution is x=5